### v3

# PRISM: Leveraging Prototype Patient Representations with Feature-Missing-Aware Calibration for EHR Data Sparsity Mitigation

Abstract
Electronic Health Record (EHR) data, while rich in information, often suffers from sparsity, posing significant challenges in predictive modeling. Traditional imputation methods inadequately distinguish between real and imputed data, leading to potential inaccuracies in models. Addressing this, we introduce PRISM, a novel approach that indirectly imputes data through prototype representations of similar patients, thus ensuring denser and more accurate embeddings. PRISM innovates further with a feature confidence learner module, which evaluates the reliability of each feature in light of missing data. Additionally, it incorporates a novel patient similarity metric that accounts for feature confidence, avoiding overreliance on imprecise imputed values. Our extensive experiments on the MIMIC-III and MIMIC-IV datasets demonstrate PRISM’s superior performance in predicting in-hospital mortality and 30-day readmission tasks, showcasing its effectiveness in handling EHR data sparsity. For the sake of reproducibility and further research, we have made the code publicly available at https://github.com/yhzhu99/PRISM.
footnotetext: $\dagger$ Corresponding author.
1 Introduction
Electronic Health Records (EHR) have become indispensable in modern healthcare, offering a rich source of data that chronicles a patient’s medical history. Over recent years, machine learning techniques have garnered significant attention in leveraging EHR data to inform clinical decisions. Such applications range from predicting the survival risk of patients Ma et al. (2020b); Zhang et al. (2021) to forecasting early mortality outcomes Yan et al. (2020); Ma et al. (2022); Gao et al. (2023); Zhu et al. (2023a).
Working with time-series EHR data presents challenges due to its inherent sparsity. Factors such as data corruption Chen and Zhang (2020), expensive examinations Ford and Ford (2000), and safety considerations Zhou et al. (2020) result in missing observations; for instance, not all indicators are captured during every patient visit Che et al. (2018). Imputed values, while necessary, are not genuine reflections of a patient’s condition and can introduce noise, diminishing model accuracy Zhang et al. (2022). Given that most machine learning models cannot process NaN (Not a Number) inputs, this sparsity necessitates imputation, complicating EHR predictive modeling. While most existing works tackling EHR sparsity have tried to perform the imputation task directly on raw data, based on modeling the health status trajectory of the whole training set, this approach is similar to that of matrix completion methods, such as MICE Van Buuren and Groothuis-Oudshoorn (2011), non-negative matrix factorization (NMF) Wang and Zhang (2012), and compressed sensing Lopes (2013). However, they fail to capture the time-series characteristics of longitudinal EHR data for each patient. Also, previous EHR-specific models with strategies of recalibrating patient representations based on attentive feature importance Ma et al. (2020a, 2023) exhibit an issue of inadvertently prioritizing imputed features, potentially introducing inaccuracies Liu et al. (2023). Thus, addressing the sparsity in time-series EHR data requires a focus on capturing relevant feature representations across visits within a patient or among patients that are essential for predictive purposes. The primary objective is to discern and highlight crucial features while diminishing the impact of irrelevant, redundant, or missing ones.
Intuitively, incorporating knowledge from similar patients as an indirect method of imputation has the potential to enhance patient representations in the context of sparse EHR data. Such a knowledge-driven approach mirrors the real-world clinical reasoning processes, harnessing patterns observed in related patient cases Suo et al. (2018). Regarding the identifying similar patient process, however, existing models also face a significant limitation: they cannot distinguish between actual and imputed data Zhang et al. (2021, 2022). Consider two patients who have the same lab test feature value. For patient A, this value originates from the actual data, while for patient B, it is an imputed value. Current similarity metrics, whether they are L1, L2 distance, cosine similarity Lee et al. (2015), handcrafted metrics Huang et al. (2019), or learning-based methods Suo et al. (2018), interpret these values identically. This results in a potentially misleading perception of similarity.
Given these insights, we are confronted with a pressing challenge: How can we effectively mitigate the sparsity issue in EHR data caused by missing recorded features, ensuring accurate and dense patient representations?
To address this, we introduce PRISM model that leverages prototype similar patients representations at the hidden state level. Unlike traditional direct imputation methods, our strategy, guided by prediction targets and with feature-missing-aware calibration, serves as a more effective imputation strategy, thus mitigating EHR data sparsity issue. Central to our approach is the feature-missing-aware calibration process in our proposed feature confidence learner module. It evaluates the reliability of each feature, considering its absence, the time since the last recorded visit, and the overall rate of missing data in the dataset. By emphasizing feature confidence, our newly designed patient similarity measure provides evaluations based not just on raw data values, but also on the varying confidence levels of each feature. Overall, our indirect imputation method distinctively differs from direct imputation methods, minimizing the introduction of misleading information.
In healthcare, the absence of data can severely compromise confidence in a prognosis. Addressing and understanding the challenges of these missing features is of utmost importance. In light of this, our work, PRISM seeks to bridge this gap. Our primary contributions are as follows:
  • Methodologically, we propose PRISM, a framework for learning prototype representations of similar patients, designed to mitigate EHR data sparsity. We design the feature confidence learner that evaluates and calibrates the reliability of each feature by examining its absence and associated confidence level. We also introduce the confidence-aware prototype patient learner with enhanced patient similarity measures that differentiates between varying feature confidence levels. Such feature-missing awareness through the whole pipeline allows for a more refined patients representation.
  • Experimentally, comprehensive experiments on two real-world datasets, MIMIC-III and MIMIC-IV, focusing on in-hospital mortality and 30-day readmission prediction tasks, reveal that PRISM significantly improves the quality of patient representations against EHR data sparsity. It outperforms the best-performing baseline model with relative improvements of 6.40% and 2.78% in mortality and 1.38% and 1.63% in readmission on AUPRC for MIMIC-III and MIMIC-IV, respectively. Further ablation studies and detailed experimental analysis underline PRISM’s effectiveness, robustness, adaptability, and efficiency.
2 Related Work
In the realm of EHR data analysis, irregular sampling often leads to significant data sparsity, presenting substantial challenges in modeling. Previous methods on sparse EHR data predominantly fall under two categories: indirect and direct imputation.
Direct Imputation
Direct imputation methods aim to estimate missing features or incorporate missing information directly. Traditional matrix imputation techniques, such as MICE Van Buuren and Groothuis-Oudshoorn (2011), non-negative matrix factorization (NMF) Wang and Zhang (2012), compressed sensing Lopes (2013), or naive zero, mean, or median imputation, rely on similar rows or columns to fill in missing data. However, these methods often operate under the assumption that patient visits are independent and features are missing at random. GRU-D Che et al. (2018) adopts a more targeted approach by introducing missing statuses into the GRU network. By utilizing time interval and missing mask information, GRU-D treats missing data as “Informative Missing”. Extending these capabilities, ConCare Ma et al. (2020b) and AICare Ma et al. (2023) first apply directly imputed EHR data, then incorporate multi-head self-attention mechanisms to refine feature embeddings. This ensures contextual relevance across diverse healthcare situations, regardless of data completeness.
Indirect Imputation
An alternative approach is evident in GRASP Zhang et al. (2021) and M3Care Zhang et al. (2022), which emphasizes the use of similar patient representations to derive meaningful information with the insight that the information observed from similar patients can be utilized as guidance for the current patient’s prognosis Zhang et al. (2021). However, accurately measuring patient similarity is intrinsically challenging, especially when features might be imputed with potentially misleading information. Many traditional works, such as  Lee et al. (2015) and  Huang et al. (2019), have resorted to fixed formulas like cosine similarity and Euclidean distance to gauge patient similarity. While these methods are straightforward, they often suffer from scalability and performance limitations. A more dynamic approach is seen in  Suo et al. (2018), which adopts metric learning with triplet loss. This technique focuses on learning the relative distances between patients, where distances have an inverse correlation with similarity scores.
However, a shared oversight across the aforementioned methods is the underestimation of the impact of missing features. This is evident both during the recalibration of features and when assessing patient similarities, as highlighted in the introduction and essential to tackle the EHR sparsity issue.
3 Problem Formulation
EHR Datasets
EHR datasets consist of a sequence of dynamic and static information for each patient. Assuming that there are $F$ features in total, $D$ dynamic features (e.g., lab tests and vital signs) and $S$ static features (e.g., sex and age), where $F=D+S$, at every clinical visit $t$. The features recorded at visit $t$ can be denoted as $\bm{x}_{t}\in\mathbb{R}^{F},t=1,2,\cdots,T$, with total $T$ visits. The dynamic feature information can be formulated as a 2-dimensional matrix $\bm{d}\in\mathbb{R}^{T\times D}$, along with static information denoted as 1-dimensional matrix $\bm{s}\in\mathbb{R}^{S}$. In addition, to differentiate between categorical and numerical variables within dynamic features, we employ one-hot encoding for categorical variables. Due to the inherent sparsity of EHR data, we incorporate feature missingness as part of our inputs. At a global view, we define the missing representation, denoted as $\rho_{i}$, to be the presence rate of the $i$-th feature within the entire dataset. From a local view, the missing representation, $\tau_{i,t}$, signifies the time interval since the last recorded visit that contains the $i$-th feature up to the $t$-th visit.
Predictive Objective
Prediction objective is presented as a binary classification task. Given each patient’s EHR data $\bm{X}=[\bm{x_{1}},\bm{x_{2}},\cdots,\bm{x_{T}}]^{\top}\in\mathbb{R}^{T\times F}$ and feature missing status $\{\rho,\tau\}$ as input, where each $\bm{x_{t}}$ consists of dynamic features and static features representation, the model attempts to predict the specific clinical outcome, denoted as $y$. The objective is formulated as $\hat{y}=\text{Model}(\bm{X},\{\rho,\tau\})$. In the MIMIC dataset’s in-hospital mortality prediction task, the goal is to predict the discharge status (00 for alive, $1$ for deceased) based on the initial 48-hour window of an ICU stay. Similarly, the 30-day readmission task predicts if a patient will be readmitted in 30 days (00 for no readmission, $1$ for readmission).
4 Methodology
4.1 Overview
Figure 1: Overall model architecture of our proposed method. “Rep.” means “Representation”.
Figure 1 shows the overall pipeline of PRISM. It consists of four main sub-modules below.
  • Feature-Isolated Embedding Module applies GRU and MLP backbone separately to dynamic features and static features. Each dynamic feature learns historical representations over multiple time steps. In order to align with the original attribute information of each feature, the features are learned in isolation from each other.
  • Feature Confidence Learner improves self-attention model by introducing the feature missing status (the global dataset-level and local patient-level missing representations of the features), collaboratively learning the confidence level of the features and the confidence-calibrated feature importance.
  • Confidence-Aware Prototype Patient Learner improves the measure of patient similarity based on patient representation and the confidence level of features learned from feature missing status. It then applies GNN to learn prototype similar patients.
  • Prototype Representation Fusion Module fuses the prototype patient embedding and the patient’s own representation, further enhancing the hidden state representation of the patient that is affected by missing data.
4.2 Feature-Isolated Embedding
In this module, static and dynamic features are learned individually, yielding feature representations of unified dimensions $f$.
Static Features Embedding Static features remain constant at each visit. Hence, we opt for a single-layer MLP to map each static feature into the feature dimensions $f$:
$\begin{array}[]{cc}\bm{h_{s_{i}}}=\mathrm{MLP}_{i}(\bm{s_{i}}),&i=1,2,\cdots,S\end{array}$ (1)
where $\bm{s_{i}}$ is the $i$-th static feature. We employ $S$ distinct-parameter MLPs for feature mappings.
Dynamic Features Embedding To ensure that each feature’s individual statistics, e.g. missing status can be incorporated with the corresponding feature, we adopt multi-channel GRU structure to avoid feature interaction at this stage. Each feature is embed with an isolated GRU:
$\begin{array}[]{cc}\bm{h_{d_{i}}}=\mathrm{GRU}_{i}(\bm{d_{i}}),&i=1,2,\cdots,D\end{array}$ (2)
where $\mathrm{GRU}_{i}$ represents the GRU network applied to the $i$-th dynamic feature $\bm{d_{i}}\in\mathbb{R}^{T\times m}$. Furthermore, $\bm{h_{d_{i}}}\in\mathbb{R}^{T\times f}$ signifies the embedding of the $i$-th dynamic feature. The in-channel of GRU is the feature recorded dimension $m$, and the out-channel is the unified $f$.
Then we employ a stack operation to integrate information from both static and dynamic features. This necessitates initially replicating the static features embeddings to each time visits: $\bm{h_{s_{i}}}\in\mathbb{R}^{f}\rightarrow\bm{h^{\prime}_{s_{i}}}\in\mathbb{R}^{T\times f}$. The stack operation is represented as follows:
$\bm{h}=\mathrm{stack}(\bm{h^{\prime}_{s_{1}}},\bm{h^{\prime}_{s_{2}}},\cdots,\bm{h^{\prime}_{s_{S}}},\bm{h_{d_{1}}},\cdots,\bm{h_{d_{D}}})$ (3)
where $\bm{h}\in\mathbb{R}^{F\times T\times f}$ represents the overall embeddings of features.
4.3 Feature Confidence Learner
Existing models often adopt various ways to enhance patient representations in order to mitigate the noise introduced by processing sparse EHR data. However, these models often use imputed data and ignore the impact of feature missing status, thus reducing the credibility of the learned hidden representation. We design a measurement called “feature confidence”, which represents the reliability of the input feature values for each patient and each visit. In addition, we have incorporated this measure into the self-attention mechanism as a recalibration module to elevate low-confidence features’ attention.
Feature Missing Status Representation We introduce two feature missing status-related variables $\rho$ and $\tau$. Global missing representation $\rho_{i}$ represents the presence rate of the $i$-th feature in the original dataset. For example, if the dataset collects a total of 100 data records during the visits of all patients, but certain feature is only recorded two times, then the $\rho$ for this feature is $\frac{2}{100}=0.02$. Local missing representation $\tau_{i,t}$ represents the time interval since the last record of this feature at the current visit. There are two special cases: case 1) If the feature is recorded at the current visit, it is marked as 0; case 2) if the feature has never been recorded before the current visit, it is marked as infinity:
$\tau_{i,t}=\begin{cases}0&\text{if case 1)}\\ \infty&\text{if case 2)}\\ t-t^{*}&\text{otherwise}\end{cases}$ (4)
where $t^{*}$ is the time of the last record of the $i$-th feature.
Missing-Aware Self-Attention To calculate the feature confidence, we comprehensively consider the missing feature status in the dataset, including the global missing representation $\rho$ and local missing representation $\tau$ , and integrate them into a self-attention mechanism module.
First, the $Query$ vector is computed from the hidden representation of the last time step $T$, while the $Key$ and $Value$ vectors are computed from the hidden representations of all time steps:
${q}_{i,T}=\bm{W}_{i}^{q}\cdot{h}_{i,T}$ (5)
${k}_{i,t}=\bm{W}_{i}^{k}\cdot{h}_{i,t}$ (6)
${v}_{i,t}=\bm{W}_{i}^{v}\cdot{h}_{i,t}$ (7)
where ${q}_{i,T}$, ${k}_{i,t}$, ${v}_{i,t}$ are the $Query$, $Key$, $Value$ vectors respectively, and $\bm{W}_{i}^{q}$, $\bm{W}_{i}^{k}$, $\bm{W}_{i}^{v}$ are the corresponding projection matrices. Following this, we compute the attention weights as follows:
$\bm{{\alpha}}_{i,*,t}={softmax}(\frac{{q}_{i,T}\bm{{k}}_{*,t}^{\top}}{\sqrt{d_{k}}})$ (8)
Subsequently, the feature confidence learner takes into account both feature missing status and attention weights to compute the feature confidence, which serves as an uncertainty reference when identifying similar patients in subsequent steps:
$C_{i,t}=\begin{cases}\tanh\left(\frac{\alpha_{i,t}}{\omega_{i,t}}\right)&\text{if }\tau_{i,t}\neq\infty\\ \beta\cdot\rho_{i}&\text{if }\tau_{i,t}=\infty\end{cases}$ (9)
where $\omega_{i,t}$ and $\alpha_{i,t}$ are defined as:
$\omega_{i,t}={\gamma_{i}\cdot\log(e+(1-\alpha_{i,t})\cdot\tau_{i,t})}$ (10)
$\alpha_{i,t}=AvgPool(\bm{\alpha}_{i,*,t})$ (11)
Here, the global missing parameter $\beta$ is a learnable parameter for global missing representation and the time-decay ratio $\gamma_{i}$ is a feature-specific learnable parameter to reflect the influence of local missing representation as time interval increases. The calculation of feature confidence is divided into two scenarios:
  • When the feature has been examined in previous visits, the authentic values of the same patient are often used to complete features. However, the feature confidence of imputed values for this feature should significantly diminish when:
    • time interval $\tau_{i,t}$ is large. As the time interval increases, the feature confidence will decay sharply.
    • the time-decay ratio $\gamma_{i}$ is high. The higher the time-decay ratio, the more severe the decay of the feature confidence level, and only the most recent recorded data matters.
  • Until the present visit, the examination of this specific feature has not been conducted. In this scenario, the imputed values are derived from other patients and the global missing representation $\rho$ is selected to depict the feature confidence for the current patient.
Finally, based on feature confidence, we can obtain the calibrated attention weights $\bm{\alpha^{*}}$ and further hidden representations $\bm{z}$.
$\bm{\alpha^{*}}=\epsilon\cdot\bm{\alpha}+(1-\epsilon)\cdot\bm{C}$ (12)
$\bm{z}=\bm{\alpha^{*}}\bm{V}$ (13)
where $\epsilon$ is a learnable parameter, $\bm{\alpha}$, $\bm{C}$, $\bm{\alpha^{*}}$, $\bm{V}$, $\bm{z}$ $\in\mathbb{R}^{T\times F}$ are attention weights, feature confidence, calibrated attention weights, value vector and learned representation of the input patient.
4.4 Confidence-Aware Prototype Patient Learner
To comprehensively account for the implications of absent features in the evaluation and identification of analogous patients, we integrate the feature confidence status into our confidence-aware module for discovering cohorts of similar patients. Based on the graph convolutional network (GCN), we compute the similarity score by factoring in the acquired feature confidence pertaining to the individual patient. This approach facilitates the preferential selection of patients who are more similar to the subject patient.
Confidence-Aware Patient Similarity Measure To find similar patients for the current patient, we first need to calculate the similarity with other patients. This is achieved through the confidence-aware patient similarity measure, resulting in a similarity score matrix as $\bm{A}$. Specifically, the similarity score $\phi_{i,j}(\bm{z}_{i},\bm{z}_{j};\bm{C}_{i},\bm{C}_{j})$ between the $i$-th and $j$-th ($i\neq j$) patients is defined as Equation 14. Note that when $i=j$, indicating the comparison of a patient with themselves, the similarity score is defined in $\phi_{i,j}=1$.
$\phi_{i,j}=\frac{1}{(1-\zeta)\cdot\frac{1}{F}(\psi_{i,j}^{(z)}(\bm{z}_{i},\bm{z}_{j})+\zeta\cdot\psi_{i,j}^{(C)}(\bm{C}_{i},\bm{C}_{j}))}$ (14)
where $\psi_{i,j}^{(z)}(\bm{z}_{i},\bm{z}_{j})$ measures the similarity of the patients’ representations, $\psi_{i,j}^{(C)}(\bm{C}_{i},\bm{C}_{j})$ measures the confidence level of the two patients in their respective feature representations, and $\zeta$ serves as learnable weight to balance the two. They are defined as follows respectively:
$\psi_{i,j}^{(z)}(\bm{z}_{i},\bm{z}_{j})=||\bm{z_{i}}-\bm{z_{j}}||_{2}^{2}$ (15)
$\psi_{i,j}^{(C)}(\bm{C}_{i},\bm{C}_{j})=\sum_{k=1}^{F}\exp(1-{{C}_{i,k}})\cdot\exp(1-{{C}_{j,k}})$ (16)
Prototype Patients Cohort Discovery To compute the enhanced representation information of similar patients, we design the prototype similar patient cohort discovery module. First, we utilize the K-Means clustering algorithm to cluster all $F$ features into $K$ groups. Then we select the center vectors of $K$ clusters as the initial $K$ prototypes $\ \bm{\mathcal{G}_{k}},\ k=1,2,\cdots,K$. With convolutions occurring within these graphs, the graph structure dynamically evolves across epochs.
Based on the aforementioned similarity measure, we can find the most similar prototype for each sample, making samples within the clusters more similar to each other during the learning process; at the same time, the prototypes also become more representative during the learning process.
After concatenating the hidden representation $\bm{z}$, computed in the previous step, with $\bm{\mathcal{G}}$, we feed this along with the similarity matrix $\bm{A}$ into a GCN. Note that the feature confidence $\mathbf{C}$ of prototype patients is initially set to $1$ for each feature. However, during training, both $\bm{\mathcal{G}}$ and the corresponding $\mathbf{C}$ are adjusted. Consequently, we use the respective $\bm{\mathcal{G}}$ for each patient for representation enhancement:
$\bm{\mathcal{G}^{*}}={\rm MLP(GCN}({\mathrm{concat}(\bm{Z},\bm{\mathcal{G}})},\bm{A}))$ (17)
where $\bm{\mathcal{G}^{*}}$ is the updated prototype representation by GCN and MLP.
4.5 Prototype Representation Fusion Module
Currently, there are two learned hidden representations, one is $\bm{z}$ obtained through the missing-aware self-attention module, and the other is $\bm{\mathcal{G}}$ obtained through the prototype similar patient cohort discovery module. A fusion module thus is needed to comprehensively fuse the information:
$\bm{z_{i}^{*}}=\eta\cdot\bm{\mathcal{G}_{i}}+(1-\eta)\cdot\bm{z_{i}}$ (18)
where $\eta$ is a learnable weight parameter, $\bm{\mathcal{G}_{i}}$ is the corresponding prototype of the $i$-th sample.
Finally, the fused representation $\bm{z^{*}}$ is expected to predict downstream tasks. We sequentially pass $\bm{z^{*}}$ through two-layer GRU and a single-layer MLP network to obtain the final prediction results $\hat{y}$:
$\hat{y}={\rm MLP(GRU}(\bm{z^{*}}))$ (19)
The BCE Loss is selected as the loss function for the binary mortality outcome prediction task:
$\mathcal{L}(\hat{y},y)=-\frac{1}{n}\sum_{i=1}^{n}(y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i}))$ (20)
where $n$ is the number of patients within one batch, $\hat{y}\in[0,1]$ is the predicted probability and $y$ is the ground truth.
Table 1: In-hospital mortality and readmission prediction results on MIMIC-III and MIMIC-IV datasets. The reported score is in the form of mean±std of 5 runs. Bold indicates the top performance. All metrics are multiplied by 100 for readability purposes.
Dataset MIMIC-III Mortality MIMIC-III Readmission MIMIC-IV Mortality MIMIC-IV Readmission
Metric AUPRC($\uparrow$) AUROC($\uparrow$) AUPRC($\uparrow$) AUROC($\uparrow$) AUPRC($\uparrow$) AUROC($\uparrow$) AUPRC($\uparrow$) AUROC($\uparrow$)
MICE 52.40±0.57 85.43±0.23 50.61±0.47 77.97±0.60 49.89±0.78 84.37±0.13 45.28±0.38 75.84±0.25
GRU-D 45.31±3.22 81.72±2.46 42.13±3.19 73.31±1.65 48.79±1.57 85.02±0.50 45.03±0.71 76.47±0.53
RETAIN 51.76±0.86 85.57±0.43 47.53±0.48 77.42±0.38 54.06±0.71 86.24±0.36 49.93±0.73 78.54±0.38
AdaCare 52.28±0.50 85.73±0.19 48.76±0.35 77.65±0.32 50.45±0.80 83.96±0.13 48.57±0.29 77.00±0.20
ConCare 51.45±0.76 86.18±0.14 47.45±0.96 77.74±0.32 49.97±1.08 85.41±0.40 47.17±0.84 77.47±0.19
GRASP 53.59±0.33 86.54±0.17 50.21±0.22 78.14±0.35 54.41±0.46 86.08±0.17 50.22±0.26 78.50±0.22
M3Care 51.68±1.03 86.23±0.42 49.00±0.71 78.00±0.55 52.95±0.71 84.90±0.37 49.22±0.69 77.31±0.42
SAFARI 45.92±1.01 85.10±0.24 45.59±0.35 77.01±0.21 46.58±0.55 46.58±0.55 44.78±0.69 76.05±0.38
AICare 51.37±0.70 85.40±0.48 47.06±1.16 76.23±0.84 49.76±0.86 84.62±0.28 45.88±1.12 76.07±0.43
PRISM-${}_{\text{-proto.}}$ 55.52±0.34 87.28±0.11 51.17±0.25 78.66±0.22 55.76±0.90 86.82±0.16 50.75±0.65 79.12±0.44
PRISM-${}_{\text{-calib.}}$ 56.16±0.42 87.33±0.22 49.13±2.11 77.87±0.88 55.18±0.77 86.57±0.20 50.62±0.70 78.66±0.45
PRISM 57.02±0.38 87.34±0.22 51.31±1.02 78.76±0.59 55.92±0.75 86.82±0.20 51.04±0.70 79.14±0.33
5 Experiments
We conduct the in-hospital mortality and 30-day readmission prediction task on MIMIC-III and MIMIC-IV datasets. The dataset preprocessing scripts, developed model’s code and all baseline models’ code are open-sourced for reproducibility.
5.1 EHR Datasets
Both MIMIC-III and MIMIC-IV datasets are split into 70% training set, 10% validation set and 20% test set with stratified shuffle split strategy based on patients’ end-stage mortality outcome. By default, both datasets employ the Last Observation Carried Forward (LOCF) imputation method Woolley et al. (2009).
MIMIC-III Johnson et al. (2016) (Medical Information Mart for Intensive Care) is a large, freely-available database comprising information such as demographics, vital sign measurements made at the bedside, laboratory test results, procedures, medications, caregiver notes, imaging reports, and mortality. The subsequent iteration, MIMIC-IV Johnson et al. (2023), stands as an evolved manifestation of the MIMIC-III database, encompassing data updates and partial table reconstructions. We extracted patient EHR data following Harutyunyan et al. (2019).
5.2 Evaluation Metrics
We assess the binary classification performance using AUROC, AUPRC and F1. Here we emphasize AUPRC as the main metric due to it is informative when dealing with highly imbalanced and skewed datasets Kim and Hwang (2022); Davis and Goadrich (2006) as MIMIC datasets.
5.3 Baseline Models
Imputation-based methods
include MICE Van Buuren and Groothuis-Oudshoorn (2011) addresses missing data in EHR through iterative imputation, with subsequent analysis using an LSTM model Hochreiter and Schmidhuber (1997). GRU-D Che et al. (2018) incoporates both the last observed and global mean values in the GRU network to handle missing values,
EHR-specific models
such as RETAIN Choi et al. (2016), AdaCare Ma et al. (2020a), ConCare Ma et al. (2020b), GRASP Zhang et al. (2021), M3Care Zhang et al. (2022), SAFARI Ma et al. (2022) and AICare Ma et al. (2023). They focus on personalized health status embeddings.
Ablation models
include PRISM-${}_{\text{-proto.}}$ that removes the confidence-aware prototype similar patient learner, and PRISM-${}_{\text{-calib.}}$ that removes the feature confidence learner for missing-feature-aware calibration process.
5.4 Experimental Results
Table 1 depicts the performance evaluation of baseline methods, PRISM, and its reduced versions for ablation study on both the MIMIC-III and MIMIC-IV datasets under two prediction tasks. Additionally, we conduct t-test based on the AUPRC metric, the PRISM’s performance improvement against all models are all statistically significant with p-value $<$ 0.01, which underscores that PRISM significantly outperforms all baseline models. Specifically, PRISM outperforms models focused solely on enhancing feature representations with attention mechanisms (like RETAIN, AdaCare, ConCare, AICare, SAFARI), by integrating missing feature status into these mechanisms for improved attention calibration and feature representation. It also exceeds models using similar patient information (such as GRASP, M3Care), showing the value of missing feature status in refining prototype patient representations for better performance. PRISM’s advantage over GRU-D, which only considers local patient visit-based feature missing status, highlights the significance of a global perspective on overall feature missing rates for effective feature representation across patients.
5.5 Ablation Study
Comparing with Reduced Versions
Compared to reduced versions, PRISM outperforms PRISM-${}_{\text{-proto.}}$ and PRISM-${}_{\text{-calib.}}$ on main metric AUPRC. This indicates that the two designed learners can enhance patient feature representations from different perspectives: the patient’s individual health data utilized by the feature confidence learner based on the attention mechanism and the prototype similar patient representations utilized by the prototype similar patient learner.
Comparing with Internal Components
To deeply explore the impact of components within each module, we conduct experiments in Table 2, showing PRISM outperforms all baselines. The symbol $\phi$ denotes similarity measure, detailed in Equation 14. The term $z$ alone indicates the use of L2 distance for patient similarity measure, whereas $z,C$ additionally incorporates feature confidence, enhancing the model’s discriminative capability. Comparing the roles of various components within the feature confidence learner, the performance when considering both global feature missing rate $\rho$ and local patient’s time interval $\tau$ is higher than considering any single component, which illustrates the necessity of considering the feature missing status from both global and local perspectives. When only considering the local perspective, its performance actually worsens, which is consistent with our observation of the performance of GRU-D. As for confidence-aware prototype patient learner, the performance of confidence-aware patient similarity measurement surpasses that without considering feature confidence, which also shows the impact of missing feature status on measuring similar patients.
Table 2: Performance comparison of internal components on the MIMIC-IV mortality prediction task. The reported score is in the form of mean±std. “Feat. Conf.” means “Feature Confidence” and “Sim. Meas.” denotes “Similarity Measure”. Bold denotes the highest performance within each components, Red denotes the highest performance among all comparisons.
Comparisons Components Metrics
+$\rho$ +$\tau$ +$\phi$ AUPRC($\uparrow$) AUROC($\uparrow$)
Feat. Conf. / / / 54.42±0.43 86.22±0.30
$\surd$ / / 55.11±0.53 86.56±0.33
/ $\surd$ / 52.67±2.52 86.43±0.55
$\surd$ $\surd$ / 55.76±0.90 86.82±0.16
Sim. Meas. / / $\bm{z}$ 55.18±0.77 86.57±0.20
/ / $\bm{z}$, $\bm{C}$ 55.25±0.75 86.60±0.19
PRISM $\surd$ $\surd$ $\bm{z}$, $\bm{C}$ 55.92±0.75 86.82±0.20
5.6 Observations and Analysis
Figure 2: AUPRC Performance across 5 Sparsity Levels. PRISM significantly outperforms other models in extreme sparse scenario.
Robustness to Data Sparsity
To assess PRISM’s performance in conditions of data scarcity, we compare it with leading models such as GRASP, MICE(LSTM), and RETAIN in the MIMIC-IV mortality prediction task. We intentionally increase the missing data rates beyond the original missing rate in MIMIC-IV. The results in Figure 2 demonstrate that PRISM excels across all conditions. Notably, in situations of extreme data scarcity, PRISM significantly outperforms the other models, highlighting its robustness in sparse data.
Sensitiveness to Cohort Size and Effectiveness of Missing-Feature-Aware Module
We conduct a detailed analysis to examine the impact of prototype patients cohort diversity and the role of a missing-feature-aware module in patient prototypes. Figure 3 shows that integrating a missing-feature-aware module consistently enhances performance across various cluster sizes, as indicated by the superior AUPRC and F1 scores. Furthermore, the relatively consistent performance across different cluster sizes demonstrates that our model is not overly sensitive to the number of clusters, highlighting its adaptability and robustness in managing various cohort sizes.
(a) AUPRC
(b) F1
Figure 3: Prototype patient cohort size (MIMIC-IV, mortality). With the missing-feature-aware module, PRISM outperforms its counterpart without. It also shows PRISM is not sensitive to the cohort size.
Variations of Similarity Measures
We conduct a comparison against standard similarity metrics, as detailed in Table 3, PRISM outperforms cosine similarity, and L1 and L2 distances in all metrics. This demonstrates the effectiveness of PRISM’s confidence-aware patient similarity measure in handling sparse EHR data.
Table 3: Different similarity measures (MIMIC-IV, mortality).
Measure AUPRC($\uparrow$) AUROC($\uparrow$) F1($\uparrow$)
Cosine 55.58±0.43 86.81±0.10 45.44±1.99
L1 55.59±0.54 86.73±0.20 45.20±1.91
L2 55.38±0.51 86.76±0.09 43.69±1.25
PRISM 55.92±0.75 86.82±0.20 45.47±2.26
Model Efficiency and Complexity
In our comparative study, we evaluate the efficiency and complexity against SOTA EHR-specific models baseline models. Table 4 highlights PRISM’s balance between a relatively low total parameter count (0.882M) and a competitive runtime (69.03s).
Table 4: Computational comparison with latest SOTA EHR-specific models on MIMIC-IV. All models employ the same 128 hidden dimension and 1024 batch size hyperparameters.
Model # Parameter Runtime (5 epochs)
ConCare (AAAI-2020) 13.019M 96.62s
GRASP (AAAI-2021) 13.153M 97.68s
M3Care (KDD-2022) 1.351M 68.29s
SAFARI (TKDE-2022) 7.228M 78.97s
AICare (Patterns-2023) 8.901M 72.97s
PRISM 0.882M 69.03s
Figure 4: Cross-Feature Attention. Without (Left) / With (Right).
Cross-Feature Attention Map
Figure 4 presents the cross-feature attention map from PRISM, contrasting average attention weights with and without feature confidence calibration process. This visualization, based on a single diagnostic record from randomly selected MIMIC-IV patients, plots Key features on the x-axis against Query features on the y-axis. Notably, PRISM reduces attention on features like capillary refill rate, fraction inspired oxygen, and height, which have high missing rates (99.64%, 93.62%, and 99.61%, respectively) in the dataset. PRISM recognizes this characteristic, thereby causing the three horizontal lines on the right side of the graph to appear distinctively whiter.
6 Conclusions
In this work, we propose PRISM, a novel prototype patient representation learning framework to address the sparsity issue of EHR data. PRISM innovatively perceives and calibrates for missing features, thereby refining patient representations via a confidence-aware prototype patient learner. Significant performance improvements and detailed experimental analysis on two real-world datasets’ in-hospital mortality and 30-day readmission prediction tasks show PRISM’s effectiveness. The work marks a crucial step towards more reliable and effective utilization of EHR data in healthcare, offering a potent solution to the prevalent issue of data sparsity in clinical decision-making.
Ethical Statement
This study, involving the analysis of Electronic Health Records (EHR) using the MIMIC dataset, is committed to upholding high ethical standards. The MIMIC dataset is a de-identified dataset, ensuring patient confidentiality and privacy. It is available through a data use agreement, underscoring our commitment to responsible data handling and usage. Our approach has been designed to minimize any potential harm and to ensure that our findings are as unbiased and fair as possible, taking into account the diverse and complex nature of medical data. We have also taken rigorous steps to ensure our research aligns with these values.
Acknowledgments
This work is supported by the National Key R&D Program of China (No. 2022ZD0116401).
Appendix A Appendix
A.1 Notation Table
Table 5: Notations symbols and their descriptions
Notations
Descriptions
$N$
Number of patient samples
$T$
Number of visits for a certain patient
$D$
Number of dynamic features
$S$
Number of static features
$F$
Number of features, $F=S+D$
$\bm{d_{it}}\in\mathbb{R}^{m}$
The $i$-th feature at the $t$-th visit, where $m$ is either the number of categories (for one-hot encoding) or 1 (for numerical lab tests)
$\bm{X}\in\mathbb{R}^{T\times F}$
Clinical visit matrix of a single patient, consisting of $T$ visits
$\bm{s},\bm{d}$
Static and dynamic feature vector of a patient
$\bm{y},\bm{\hat{y}}$
Ground truth labels and prediction results
$\bm{h_{i}}\in\mathbb{R}^{T\times f},\bm{h}$
Representation of $i$-th feature learned by GRU (for dynamic features) or MLP (for static features), stacked up to formulate the representation matrix $\bm{h}$, $f$ is each feature’s embedding dimension
$\rho_{i}$
Feature presence rate of feature $i$ in the training set
$\tau_{i,t}$
Time interval from the last recorded visit of $i$-th feature at $t$-th visit
$\bm{C}\in\mathbb{R}^{T\times F}$
Learned feature confidence matrix of a patient
$\bm{z_{i}}$
$\bm{z_{i}}$ is the learned representation of the $i$-th patient after the feature calibration layer
$\bm{\alpha},\bm{\alpha^{*}}$
Learned attention weights and calibrated attention weights after the feature calibration layer
$\phi(\cdot,\cdot)$
Patient similarity measure function
$\bm{A}=(a_{i,j})$
Adjacency matrix of patients, composed of similarity score between $i$-th and $j$-th patient
$\bm{\mathcal{G}_{k}}$
Learned prototype patient representation of the $k$-th group
$\bm{z^{*}_{i}}$
Learned representation of the $i$-th patient after representation fusion layer
$\bm{W}_{\square}$
Parameter matrices of linear layers. Footnote $\square$ denotes the name of the layer
$K$
Number of similar patient groups
A.2 Feature Decay Rates Observation
Figure 5: Feature Decay Rates
Figure 5 illustrates the decay rates of adaptive learning for different features, describing how the significance of feature values diminishes over time. Elevated decay rates accentuate the importance of short-term feature patterns, whereas reduced decay rates indicate the significance of long-term persistent feature patterns. The figure reveals that our model places higher attention on the short-term dynamics of certain features, including height, heart rate and pH. The heightened decay rate enables our model to swiftly detect short-term patterns or variations in these attributes, capturing shifts in the patient’s physical state. This alignment with medical intuition aligns these features with rapid changes, often indicative of acute cases, such as shock, bleeding, infection, etc. On the other hand, attributes such as sex and systolic blood pressure values require attention from a long-term perspective.
A.3 Ablation Models Details
Ablation models include PRISM-${}_{\text{-proto.}}$ and PRISM-${}_{\text{-calib.}}$ in the “Comparing with Reduced Versions” experiment. PRISM-${}_{\text{-proto.}}$ removes the confidence-aware prototype similar patient learner and reserves the feature-missing-aware calibration process. We apply the patients hidden representation $\bm{z}$ for downstream GRU module. PRISM-${}_{\text{-calib.}}$ that removes the feature confidence learner for missing-feature-aware calibration process. As the confidence-aware protype patient learner require the feature confidence $\bm{C}$ as input, we set the $\bm{C}=1$ for each feature.
A.4 Implementation Details
We run each model 5 times using 5 different random seeds: ${0,1,2,3,4}$. All runs are trained on a single Nvidia RTX 3090 GPU with CUDA 11.8. The server’s RAM size is 64GB. We implement the model in Python 3.11.4, using PyTorch 2.0.1 Paszke et al. [2019], PyTorch Lightning 2.0.5 Falcon [2019], and pyehr Zhu et al. [2023b]. AdamW Loshchilov and Hutter [2017] is employed with a batch size of 1024 patients. All models are trained for 50 epochs, and an early stop strategy, monitored by AUPRC, is applied after 10 epochs without improvement. The learning rate ${0.01,0.001,0.0001}$ and hidden dimensions ${64,128}$ are tuned using a grid search strategy on the validation set. The searched hyperparameter for PRISM is: 128 hidden dimensions, 0.001 learning rate, and 256 prototype patient cluster numbers.