### v4

# PRISM: Leveraging Prototype Patient Representations with Feature-Missing-Aware Calibration for EHR Data Sparsity Mitigation

Abstract
Electronic Health Record (EHR) data, while rich in information, often suffers from sparsity, posing significant challenges in predictive modeling. Traditional imputation methods inadequately distinguish between real and imputed data, leading to potential inaccuracies in models. Addressing this, we introduce PRISM, a framework that indirectly imputes data through prototype representations of similar patients, thus ensuring denser and more accurate embeddings. PRISM also includes a feature confidence learner module, which evaluates the reliability of each feature in light of missing data. Additionally, it incorporates a new patient similarity metric that accounts for feature confidence, avoiding overreliance on imprecise imputed values. Our extensive experiments on the MIMIC-III, MIMIC-IV, PhysioNet Challenge 2012, eICU datasets demonstrate PRISM’s superior performance in predicting in-hospital mortality and 30-day readmission tasks, showcasing its effectiveness in handling EHR data sparsity. For the sake of reproducibility and further research, we have made the code publicly available at https://github.com/yhzhu99/PRISM.
footnotetext: Corresponding author.
1 Introduction
Electronic Health Records (EHR) have become indispensable in modern healthcare, offering a rich source of data that chronicles a patient’s medical history. Over recent years, machine learning techniques have garnered significant attention in leveraging EHR data to inform clinical decisions. Such applications range from predicting the survival risk of patients [1, 2] to forecasting early mortality outcomes [3, 4, 5, 6].
Working with time-series EHR data presents challenges due to its inherent sparsity. Factors such as data corruption [7], expensive examinations [8], and safety considerations [9] result in missing observations; for instance, not all indicators are captured during every patient visit [10]. Imputed values, while necessary, are not genuine reflections of a patient’s condition and can introduce noise, diminishing model accuracy [11]. Given that most machine learning models cannot process NaN (Not a Number) inputs, this sparsity necessitates imputation, complicating EHR predictive modeling. While most existing works tackling EHR sparsity have tried to perform the imputation task directly on raw data, based on modeling the health status trajectory of the whole training set, this approach is similar to that of matrix completion methods, such as MICE [12], non-negative matrix factorization (NMF) [13], and compressed sensing [14]. However, they fail to capture the time-series characteristics of longitudinal EHR data for each patient. Also, previous EHR-specific models with strategies of recalibrating patient representations based on attentive feature importance [15, 16] exhibit an issue of inadvertently prioritizing imputed features, potentially introducing inaccuracies [17]. Thus, addressing the sparsity in time-series EHR data requires a focus on capturing relevant feature representations across visits within a patient or among patients that are essential for predictive purposes. The primary objective is to discern and highlight crucial features while diminishing the impact of irrelevant, redundant, or missing ones.
Intuitively, incorporating knowledge from similar patients as an indirect method of imputation has the potential to enhance patient representations in the context of sparse EHR data. Such a knowledge-driven approach mirrors the real-world clinical reasoning processes, harnessing patterns observed in related patient cases [18]. Regarding the identifying similar patient process, however, existing models also face a significant limitation: they cannot distinguish between actual and imputed data [2, 11]. Consider two patients who have the same lab test feature value. For patient A, this value originates from the actual data, while for patient B, it is an imputed value. Current similarity metrics, whether they are L1, L2 distance, cosine similarity [19], handcrafted metrics [20], or learning-based methods [18], interpret these values identically. This results in a potentially misleading perception of similarity.
Given these insights, we are confronted with a pressing challenge: How can we effectively mitigate the sparsity issue in EHR data caused by missing recorded features, ensuring accurate and dense patient representations?
To address this, we introduce PRISM model that leverages prototype similar patients representations at the hidden state level. Unlike traditional direct imputation methods, our strategy, guided by prediction targets and with feature-missing-aware calibration, serves as a more effective imputation strategy, thus mitigating EHR data sparsity issue. Central to our approach is the feature-missing-aware calibration process in our proposed feature confidence learner module. It evaluates the reliability of each feature, considering its absence, the time since the last recorded visit, and the overall rate of missing data in the dataset. By emphasizing feature confidence, our newly designed patient similarity measure provides evaluations based not just on raw data values, but also on the varying confidence levels of each feature. Overall, our indirect imputation method distinctively differs from direct imputation methods, minimizing the introduction of misleading information.
In healthcare, the absence of data can severely compromise confidence in a prognosis. Addressing and understanding the challenges of these missing features is of utmost importance. In light of this, PRISM seeks to bridge this gap. Our primary contributions are:
  • Methodologically, we propose PRISM, a framework for learning prototype representations of similar patients, designed to mitigate EHR data sparsity. We design the feature confidence learner that evaluates and calibrates the reliability of each feature by examining its absence and associated confidence level. We also introduce the confidence-aware prototype patient learner with enhanced patient similarity measures that differentiates between varying feature confidence levels. Compared to existing SOTA baselines, i.e. GRASP [2] and M3Care [11], PRISM provides a refined feature calibration, and further missing-aware similarity measure helps to identify more related patient representation, with missing feature status elaborately taking into account.
  • Experimentally, comprehensive experiments on four real-world datasets, MIMIC-III, MIMIC-IV, Challenge-2012, and eICU, focusing on in-hospital mortality and 30-day readmission prediction tasks, reveal that PRISM significantly improves the quality of patient representations against EHR data sparsity. PRISM outperforms the best-performing baseline model with relative improvements of 6.40%, 2.78%, 1.51% and 11.01% in mortality on AUPRC for four datasets. In terms of readmission task, PRISM obtains relative improvements of 1.38% and 1.63% on AUPRC for MIMIC-III and MIMIC-IV, respectively. Further ablation studies and detailed experimental analysis underline PRISM’s effectiveness, robustness, adaptability, and efficiency.
2 Related Work
In the realm of EHR data analysis, irregular sampling often leads to significant data sparsity, presenting substantial challenges in modeling. Previous methods on sparse EHR data predominantly fall under two categories: direct imputation in the raw data space and indirect imputation within the feature representation space.
Direct Imputation
Direct imputation methods aim to estimate missing features or incorporate missing information directly. Traditional matrix imputation techniques, such as MICE [12], non-negative matrix factorization (NMF) [13], compressed sensing [14], or naive zero, mean, or median imputation, rely on similar rows or columns to fill in missing data. However, these methods often operate under the assumption that patient visits are independent and features are missing at random. GRU-D [10] adopts a more targeted approach by introducing missing statuses into the GRU network. By utilizing time interval and missing mask information, GRU-D treats missing data as “Informative Missing”. Extending these capabilities, ConCare [1] and AICare [16] first apply directly imputed EHR data, then incorporate multi-head self-attention mechanisms to refine feature embeddings. This ensures contextual relevance across diverse healthcare situations, regardless of data completeness.
Indirect Imputation
As demonstrated in GRASP [2] and M3Care [11], they emphasize the use of similar patient representations to derive meaningful information with the insight that the information observed from similar patients can be utilized as guidance for the current patient’s prognosis [2]. However, accurately measuring patient similarity is intrinsically challenging, especially when features might be imputed with potentially misleading information. Many traditional works, such as  [19] and  [20], have resorted to fixed formulas like cosine similarity and Euclidean distance to gauge patient similarity. While these methods are straightforward, they often suffer from scalability and performance limitations. A more dynamic approach is seen in  [18], which adopts metric learning with triplet loss. This technique focuses on learning the relative distances between patients, where distances have an inverse correlation with similarity scores.
However, a shared oversight across the aforementioned methods is the underestimation of the impact of missing features. This is evident both during the recalibration of features and when assessing patient similarities, as highlighted in the introduction and essential to tackle the EHR sparsity issue.
3 Problem Formulation
Table 1: Notations symbols and their descriptions
Notations Descriptions
$N$ Number of patient samples
$T$ Number of visits for a certain patient
$D$ Number of dynamic features
$S$ Number of static features
$F$ Number of features, $F=S+D$
$\bm{d_{it}}\in\mathbb{R}^{m}$ The $i$-th feature at the $t$-th visit, where $m$ is either the number of categories (for one-hot encoding) or 1 (for numerical lab tests)
$\bm{X}\in\mathbb{R}^{T\times F}$ Clinical visit matrix of a single patient, consisting of $T$ visits
$\bm{s},\bm{d}$ Static and dynamic feature vector of a patient
$\bm{y},\bm{\hat{y}}$ Ground truth labels and prediction results
$\bm{h_{i}}\in\mathbb{R}^{T\times f},\bm{h}$ Representation of $i$-th feature learned by GRU (for dynamic features) or MLP (for static features), stacked to form the representation matrix $\bm{h}$, $f$ is each feature’s embedding dimension
$\rho_{i}$ Feature presence rate of feature $i$ in training set
$\tau_{i,t}$ Time interval from the last recorded visit of $i$-th feature at $t$-th visit
$\bm{C}\in\mathbb{R}^{T\times F}$ Learned feature confidence matrix of a patient
$\bm{z_{i}}$ $\bm{z_{i}}$ is the learned representation of the $i$-th patient after the feature calibration layer
$\bm{\alpha},\bm{\alpha^{*}}$ Learned attention weights and calibrated attention weights after the feature calibration layer
$\phi(\cdot,\cdot)$ Patient similarity measure function
$\bm{A}=(a_{i,j})$ Adjacency matrix of patients, composed of similarity score between $i$-th and $j$-th patient
$\bm{\mathcal{G}_{k}}$ Learned prototype patient representation of the $k$-th group
$\bm{z^{*}_{i}}$ Learned representation of the $i$-th patient after representation fusion layer
$\bm{W}_{\square}$ Parameter matrices of linear layers. Footnote $\square$ denotes the name of the layer
$K$ Number of similar patient groups
EHR Datasets
EHR datasets consist of a sequence of dynamic and static information for each patient. Assuming that there are $F$ features in total, $D$ dynamic features (e.g., lab tests and vital signs) and $S$ static features (e.g., sex and age), where $F=D+S$, at every clinical visit $t$. The features recorded at visit $t$ can be denoted as $\bm{x}_{t}\in\mathbb{R}^{F},t=1,2,\cdots,T$, with total $T$ visits. The dynamic feature information can be formulated as a 2-dimensional matrix $\bm{d}\in\mathbb{R}^{T\times D}$, along with static information denoted as 1-dimensional matrix $\bm{s}\in\mathbb{R}^{S}$. In addition, to differentiate between categorical and numerical variables within dynamic features, we employ one-hot encoding for categorical variables. Due to the inherent sparsity of EHR data, we incorporate feature missingness as inputs. At a global view, we define the missing representation, denoted as $\rho_{i}$, to be the presence rate of the $i$-th feature within the entire dataset. From a local view, the missing representation, $\tau_{i,t}$, signifies the time interval since the last recorded visit that contains the $i$-th feature up to the $t$-th visit.
Predictive Objective
Prediction objective is presented as a binary classification task. Given each patient’s EHR data $\bm{X}=[\bm{x_{1}},\bm{x_{2}},\cdots,\bm{x_{T}}]^{\top}\in\mathbb{R}^{T\times F}$ and feature missing status $\{\rho,\tau\}$ as input, where each $\bm{x_{t}}$ consists of dynamic features and static features representation, the model attempts to predict the specific clinical outcome, denoted as $y$. The objective is formulated as $\hat{y}=\text{Model}(\bm{X},\{\rho,\tau\})$. For the in-hospital mortality prediction task, the goal is to predict the discharge status (00 for alive, $1$ for deceased) based on the initial 48-hour window of an ICU stay. Similarly, the 30-day readmission task predicts if a patient will be readmitted in 30 days (00 for no readmission, $1$ for readmission).
Notation Table
Table 1 contains notation symbols and their descriptions used in the paper.
4 Methodology
4.1 Overview
Figure 1: Overall model architecture of our proposed method PRISM. “Rep.” means “Representation”.
Figure 1 shows the overall pipeline of PRISM. It consists of four main sub-modules below.
  • Feature-Isolated Embedding Module applies GRU and MLP backbone separately to dynamic features and static features. Each dynamic feature learns historical representations over multiple time steps. To align with the original attribute information of each feature, the features are learned in isolation from each other.
  • Feature Confidence Learner improves self-attention model by introducing the feature missing status (the global dataset-level and local patient-level missing representations of the features), collaboratively learning the confidence level of the features and the confidence-calibrated feature importance.
  • Confidence-Aware Prototype Patient Learner improves the measure of patient similarity based on patient representation and the confidence level of features learned from feature missing status. It then applies the graph neural network to learn prototype patients.
  • Prototype Representation Fusion Module fuses the prototype patient embedding and the patient’s own representation, further enhancing the hidden state representation of the patient that is affected by missing data.
4.2 Feature-Isolated Embedding
In this module, static and dynamic features are learned individually via Multilayer Perceptron (MLP) and Gated Recurrent Unit (GRU) networks, yielding feature representations of unified dimensions $f$.
Static Features Embedding Static features remain constant at each visit. Hence, we opt for a single-layer MLP for simplicity to map each static feature into the feature dimensions $f$:
$\begin{array}[]{cc}\bm{h_{s_{i}}}=\mathrm{MLP}_{i}(\bm{s_{i}}),&i=1,2,\cdots,S\end{array}$ (1)
where $\bm{s_{i}}$ is the $i$-th static feature. We employ $S$ distinct-parameter MLPs for feature mappings.
Dynamic Features Embedding To ensure that each feature’s individual statistics, e.g. missing status can be incorporated with the corresponding feature, we adopt multi-channel GRU structure to avoid feature interaction at this stage. Each feature is embed with an isolated GRU, a time-series model that has a proven track record of consistent performance in EHR modeling [5]:
$\begin{array}[]{cc}\bm{h_{d_{i}}}=\mathrm{GRU}_{i}(\bm{d_{i}}),&i=1,2,\cdots,D\end{array}$ (2)
where $\mathrm{GRU}_{i}$ represents the GRU network applied to the $i$-th dynamic feature $\bm{d_{i}}\in\mathbb{R}^{T\times m}$. Furthermore, $\bm{h_{d_{i}}}\in\mathbb{R}^{T\times f}$ signifies the embedding of the $i$-th dynamic feature. The in-channel of GRU is the feature recorded dimension $m$, and the out-channel is the unified $f$.
Then we employ a stack operation to integrate information from both static and dynamic features. This necessitates initially replicating the static features embeddings to each time visits: $\bm{h_{s_{i}}}\in\mathbb{R}^{f}\rightarrow\bm{h^{\prime}_{s_{i}}}\in\mathbb{R}^{T\times f}$. The stack operation is represented as follows:
$\bm{h}=\mathrm{stack}(\bm{h^{\prime}_{s_{1}}},\bm{h^{\prime}_{s_{2}}},\cdots,\bm{h^{\prime}_{s_{S}}},\bm{h_{d_{1}}},\cdots,\bm{h_{d_{D}}})$ (3)
where $\bm{h}\in\mathbb{R}^{F\times T\times f}$ represents the overall embeddings of features.
4.3 Feature Confidence Learner
Existing models adopt various ways to enhance patient representations to mitigate the noise introduced by processing sparse EHR data. However, these models often use imputed data and ignore the impact of feature missing status, thus reducing the credibility of the learned hidden representation. We design a measurement called “feature confidence”, which represents the reliability of the input feature values for each patient and each visit. In addition, we have incorporated this measure into the self-attention mechanism as a recalibration module to elevate low-confidence features’ attention.
Feature Missing Status Representation We introduce two feature missing status-related variables $\rho$ and $\tau$. Global missing representation $\rho_{i}$ represents the presence rate of the $i$-th feature in the original dataset. For example, if the dataset collects a total of 100 data records during the visits of all patients, but certain feature is only recorded two times, then the $\rho$ for this feature is $\frac{2}{100}=0.02$. Local missing representation $\tau_{i,t}$ represents the time interval since the last record of this feature at the current visit. There are two special cases: case 1) If the feature is recorded at the current visit, it is marked as 0; case 2) if the feature has never been recorded before the current visit, it is marked as infinity:
$\tau_{i,t}=\begin{cases}0&\text{if case 1)}\\ \infty&\text{if case 2)}\\ t-t^{*}&\text{otherwise}\end{cases}$ (4)
where $t^{*}$ is the time of the last record of the $i$-th feature.
Missing-Aware Self-Attention To calculate the feature confidence, we comprehensively consider the missing feature status in the dataset, including the global missing representation $\rho$ and local missing representation $\tau$, and integrate them into a self-attention mechanism module.
First, the $Query$ vector is computed from the hidden representation of the last time step $T$, while the $Key$ and $Value$ vectors are computed from the hidden representations of all time steps:
${q}_{i,T}=\bm{W}_{i}^{q}\cdot{h}_{i,T}$ (5)
${k}_{i,t}=\bm{W}_{i}^{k}\cdot{h}_{i,t}$ (6)
${v}_{i,t}=\bm{W}_{i}^{v}\cdot{h}_{i,t}$ (7)
where ${q}_{i,T}$, ${k}_{i,t}$, ${v}_{i,t}$ are the $Query$, $Key$, $Value$ vectors respectively, and $\bm{W}_{i}^{q}$, $\bm{W}_{i}^{k}$, $\bm{W}_{i}^{v}$ are the corresponding projection matrices. Following this, we compute the attention weights as follows:
$\bm{{\alpha}}_{i,*,t}={softmax}(\frac{{q}_{i,T}\bm{{k}}_{*,t}^{\top}}{\sqrt{d_{k}}})$ (8)
Subsequently, the feature confidence learner takes into account both feature missing status and attention weights to compute the feature confidence, which serves as an uncertainty reference when identifying similar patients in subsequent steps:
$C_{i,t}=\begin{cases}\tanh\left(\frac{\alpha_{i,t}}{\omega_{i,t}}\right)&\text{if }\tau_{i,t}\neq\infty\\ \beta\cdot\rho_{i}&\text{if }\tau_{i,t}=\infty\end{cases}$ (9)
where $\omega_{i,t}$ and $\alpha_{i,t}$ are defined as:
$\omega_{i,t}={\gamma_{i}\cdot\log(e+(1-\alpha_{i,t})\cdot\tau_{i,t})}$ (10)
$\alpha_{i,t}=AvgPool(\bm{\alpha}_{i,*,t})$ (11)
Here, the global missing parameter $\beta$ is a learnable parameter for global missing representation and the time-decay ratio $\gamma_{i}$ is a feature-specific learnable parameter to reflect the influence of local missing representation as time interval increases. The calculation of feature confidence is divided into two scenarios:
  • When the feature has been examined in previous visits, the authentic values of the same patient are often used to complete features. However, the feature confidence of imputed values for this feature should significantly diminish when:
    • time interval $\tau_{i,t}$ is large. As the time interval increases, the feature confidence will decay sharply.
    • the time-decay ratio $\gamma_{i}$ is high. The higher the time-decay ratio, the more severe the decay of the feature confidence level, and only the most recent recorded data matters.
  • Until the present visit, the examination of this specific feature has not been conducted. In this scenario, the imputed values are derived from other patients and the global missing representation $\rho$ is selected to depict the feature confidence for the current patient.
Finally, based on feature confidence, we can obtain the calibrated attention weights $\bm{\alpha^{*}}$ and further hidden representations $\bm{z}$.
$\bm{\alpha^{*}}=\epsilon\cdot\bm{\alpha}+(1-\epsilon)\cdot\bm{C}$ (12)
$\bm{z}=\bm{\alpha^{*}}\bm{V}$ (13)
where $\epsilon$ is a learnable parameter, $\bm{\alpha}$, $\bm{C}$, $\bm{\alpha^{*}}$, $\bm{V}$, $\bm{z}$ $\in\mathbb{R}^{T\times F}$ are attention weights, feature confidence, calibrated attention weights, value vector and learned representation of the input patient.
4.4 Confidence-Aware Prototype Patient Learner
To comprehensively account for the implications of absent features in the evaluation and identification of analogous patients, we integrate the feature confidence status into our confidence-aware module for discovering cohorts of similar patients. Inspired by GRASP [2]’s graph convolutional network (GCN) framework in EHR modeling, we further compute and integrate the similarity score within the GCN by factoring in the acquired feature confidence pertaining to the individual patient, aiming to facilitate the preferential selection of patients who are more similar to the subject patient.
Confidence-Aware Patient Similarity Measure To find similar patients for the current patient, we first need to calculate the similarity with other patients. This is achieved through the confidence-aware patient similarity measure, resulting in a similarity score matrix as $\bm{A}$. Specifically, the similarity score $\phi_{i,j}(\bm{z}_{i},\bm{z}_{j};\bm{C}_{i},\bm{C}_{j})$ between the $i$-th and $j$-th ($i\neq j$) patients is defined as Equation 14. Note that when $i=j$, indicating the comparison of a patient with themselves, the similarity score is defined in $\phi_{i,j}=1$.
$\phi_{i,j}=\frac{1}{(1-\zeta)\cdot\frac{1}{F}(\psi_{i,j}^{(z)}(\bm{z}_{i},\bm{z}_{j})+\zeta\cdot\psi_{i,j}^{(C)}(\bm{C}_{i},\bm{C}_{j}))}$ (14)
where $\psi_{i,j}^{(z)}(\bm{z}_{i},\bm{z}_{j})$ measures the similarity of the patients’ representations, $\psi_{i,j}^{(C)}(\bm{C}_{i},\bm{C}_{j})$ measures the confidence level of the two patients in their respective feature representations, and $\zeta$ serves as learnable weight to balance the two. They are defined as follows respectively:
$\psi_{i,j}^{(z)}(\bm{z}_{i},\bm{z}_{j})=||\bm{z_{i}}-\bm{z_{j}}||_{2}^{2}$ (15)
$\psi_{i,j}^{(C)}(\bm{C}_{i},\bm{C}_{j})=\sum_{k=1}^{F}\exp(1-{{C}_{i,k}})\cdot\exp(1-{{C}_{j,k}})$ (16)
Prototype Patients Cohort Discovery To compute the enhanced representation information of similar patients, we design the prototype similar patient cohort discovery module by incorporating GCN’s capability to learn relationships between graph nodes. Initially, we utilize the K-Means clustering algorithm to cluster raw patient representation ($F$ recorded or imputed features) into $K$ groups. Subsequently, we identify the center vectors of $K$ clusters as the initial $K$ prototypes $\bm{\mathcal{G}_{k}},k=1,2,\cdots,K$. These prototypes vectors, along with the patient groups, form the nodes of a graph. We then calculate the edge weights for this graph using a predefined similarity measure, resulting in an adjacency matrix $\bm{A}$. As GCN learns across epochs, the graph structure dynamically evolves. Note that the feature confidence $\mathbf{C}$ of prototype patients is initially set to $1$ for each feature. During the training phase, both $\bm{\mathcal{G}}$ and the corresponding $\mathbf{C}$ are adaptively adjusted. Consequently, the most similar and representative prototypes $\bm{\mathcal{G}}$ are identified, and samples within the clusters become more similar to each other. The process is illustrated as:
$\bm{\mathcal{G}^{*}}={\rm MLP(GCN}({\mathrm{concat}(\bm{z},\bm{\mathcal{G}})},\bm{A}))$ (17)
where $\bm{\mathcal{G}^{*}}$ is the updated prototype representation by GCN and MLP.
4.5 Prototype Representation Fusion Module
Currently, there are two learned hidden representations, one is $\bm{z}$ obtained through the missing-aware self-attention module, and the other is $\bm{\mathcal{G}}$ obtained through the prototype similar patient cohort discovery module. Thus, the patient representation is fused as:
$\bm{z_{i}^{*}}=\eta\cdot\bm{\mathcal{G}_{i}}+(1-\eta)\cdot\bm{z_{i}}$ (18)
where $\eta$ is a learnable weight parameter, $\bm{\mathcal{G}_{i}}$ is the corresponding prototype of the $i$-th sample.
Finally, the fused representation $\bm{z^{*}}$ is expected to predict downstream tasks. We sequentially pass $\bm{z^{*}}$ through two-layer GRU and a single-layer MLP network to obtain the final prediction results $\hat{y}$:
$\hat{\bm{y}}={\rm MLP(GRU}(\bm{z^{*}}))$ (19)
The BCE Loss is selected as the loss function for the binary mortality outcome prediction task:
$\mathcal{L}(\hat{y},y)=-\frac{1}{n}\sum_{i=1}^{n}(y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i}))$ (20)
where $n$ is the number of patients within one batch, $\hat{y}\in[0,1]$ is the predicted probability and $y$ is the ground truth.
5 Experimental Setups
5.1 Benchmarked Real-World Datasets
We employ MIMIC-III, MIMIC-IV, PhysioNet Challenge 2012, and eICU datasets datasets for benchmarking. All 4 datasets are split into 70% training set, 10% validation set and 20% test set with stratified shuffle split strategy based on patients’ end-stage mortality outcome. By default, we use the Last Observation Carried Forward (LOCF) imputation method [21].
  1. 1.
    MIMIC-III [22] (Medical Information Mart for Intensive Care) is a large, freely-available database comprising information such as demographics, vital sign measurements made at the bedside, laboratory test results, procedures, medications, caregiver notes, imaging reports, and mortality.
  2. 2.
    MIMIC-IV [23] as MIMIC-III dataset’s subsequent iteration, stands as an evolved manifestation of the MIMIC-III database, encompassing data updates and partial table reconstructions. We extracted patient EHR data following [24].
  3. 3.
    PhysioNet Challenge 2012 [25] (Challenge-2012) covers records from 12,000 adult ICU stays. The challenge is designed to promote the development of effective algorithms for predicting in-hospital mortality based on data from the first 48 hours of ICU admission. We utilize 35 lab test features and 5 demographic features in Challenge-2012 dataset.
  4. 4.
    eICU [26] is a large-scale, multi-center ICU database derived from over 200,000 ICU admissions across the United States between 2014 and 2015. 12 lab test features and 2 demographic features are adopted.
The statistics of datasets is in Table 2.
Table 2: Statistics of datasets after preprocessing. The proportion demonstrates the percentage of the label with value $1$. $Out.$ denotes Mortality Outcome, $Re.$ denotes Readmission.
Dataset MIMIC-III MIMIC-IV Challenge-2012 eICU
# Samples 41517 56888 4000 73386
Missing 69.87% 74.70% 84.68% 42.61%
$\text{Label}_{Out.}$ 10.62% 9.55% 13.85% 8.32%
$\text{Label}_{Re.}$ 14.74% 13.85% / /
5.2 Evaluation Metrics
We assess the binary classification performance using AUROC, AUPRC and F1. Here we emphasize AUPRC as the main metric due to it is informative when dealing with highly imbalanced and skewed datasets [27, 28] as shown in our selected datasets.
5.3 Baseline Models
We include imputation-based methods, EHR-specific models, and PRISM model with reduced modules as baseline models.
Imputation-based methods
include MICE [12] addresses missing data in EHR through iterative imputation, with subsequent analysis using an LSTM model [29]. GRU-D [10] incoporates both the last observed and global mean values in the GRU network. Additionally, GRU-D utilizes an exponential decay mechanism to manage the temporal dynamics of missing values.
EHR-specific models
such as RETAIN [30], AdaCare [15], ConCare [1], GRASP [2], M3Care [11], SAFARI [4] and AICare [16]. They focus on personalized health status embeddings.
Ablation models
include PRISM-${}_{\text{-proto.}}$ and PRISM-${}_{\text{-calib.}}$. Specifically, PRISM-${}_{\text{-proto.}}$ removes the confidence-aware prototype similar patient learner and reserves the feature-missing-aware calibration process. We apply the patients hidden representation $\bm{z}$ for downstream GRU module. PRISM-${}_{\text{-calib.}}$ removes the feature confidence learner from the missing-feature-aware calibration process. As the confidence-aware prototype patient learner requires the feature confidence $\bm{C}$ as input, we set the $\bm{C}=1$ for each feature.
5.4 Implementation Details
All runs are trained on a single Nvidia RTX 3090 GPU with CUDA 11.8. The server’s RAM size is 64GB. We implement the model in Python 3.11.4, PyTorch 2.0.1 [31], PyTorch Lightning 2.0.5 [32], and pyehr [33]. AdamW [34] is employed with a batch size of 1024 patients. All models are trained for 50 epochs with an early stopping strategy based on AUPRC after 10 epochs without improvement. The learning rate ${0.01,0.001,0.0001}$ and hidden dimensions ${64,128}$ are tuned using a grid search strategy on the validation set. The searched hyperparameter for PRISM is: 128 hidden dimensions, 0.001 learning rate, and 256 prototype patient cluster numbers. Performance is reported in the form of mean±std of 5 runs with random seeds ${0,1,2,3,4}$ for MIMIC-III and MIMIC-IV datasets, and apply bootstrapping on all test set samples 10 times for the Challenge-2012 and eICU datasets, following practices in [16]. All metric scores are multiplied by 100 for readability purposes.
6 Experimental Results and Analysis
We conduct the in-hospital mortality and 30-day readmission prediction task on MIMIC-III and MIMIC-IV datasets, in-hospital mortality prediction task on Challenge-2012 and eICU datasets. The dataset preprocessing scripts, developed model’s code and all baseline models’ code are open-sourced for reproducibility.
Table 3: Benchmarking results on MIMIC-III, MIMIC-IV, Challenge-2012, and eICU datasets. Bold indicates the best performance.
Dataset MIMIC-III Mortality MIMIC-III Readmission MIMIC-IV Mortality MIMIC-IV Readmission Challenge-2012 Mortality eICU Mortality
Metric AUPRC($\uparrow$) AUROC($\uparrow$) AUPRC($\uparrow$) AUROC($\uparrow$) AUPRC($\uparrow$) AUROC($\uparrow$) AUPRC($\uparrow$) AUROC($\uparrow$) AUPRC($\uparrow$) AUROC($\uparrow$) AUPRC($\uparrow$) AUROC($\uparrow$)
MICE 52.40±0.57 85.43±0.23 50.61±0.47 77.97±0.60 49.89±0.78 84.37±0.13 75.84±0.25 45.28±0.38 32.26±2.85 72.41±1.46 44.44±4.16 85.19±2.97
GRU-D 45.31±3.22 81.72±2.46 42.13±3.19 73.31±1.65 48.79±1.57 85.02±0.50 76.47±0.53 45.03±0.71 25.43±3.27 63.75±1.76 42.59±3.33 83.85±2.31
RETAIN 51.76±0.86 85.57±0.43 47.53±0.48 77.42±0.38 54.06±0.71 86.24±0.36 78.54±0.38 49.93±0.73 30.23±2.24 69.82±1.89 39.89±3.08 82.53±2.45
AdaCare 52.28±0.50 85.73±0.19 48.76±0.35 77.65±0.32 50.45±0.80 83.96±0.13 77.00±0.20 48.57±0.29 33.10±3.42 69.66±1.57 42.48±3.61 83.91±2.32
ConCare 51.45±0.76 86.18±0.14 47.45±0.96 77.74±0.32 49.97±1.08 85.41±0.40 77.47±0.19 47.17±0.84 30.24±3.00 70.19±2.54 44.40±4.35 85.05±2.93
GRASP 53.59±0.33 86.54±0.17 50.21±0.22 78.14±0.35 54.41±0.46 86.08±0.17 78.50±0.22 50.22±0.26 26.03±3.43 67.14±2.61 45.41±4.05 85.69±2.38
M3Care 51.68±1.03 86.23±0.42 49.00±0.71 78.00±0.55 52.95±0.71 84.90±0.37 77.31±0.42 49.22±0.69 32.63±2.54 73.26±1.67 44.95±4.32 85.44±2.56
SAFARI 45.92±1.01 85.10±0.24 45.59±0.35 77.01±0.21 46.58±0.55 46.58±0.55 76.05±0.38 44.78±0.69 28.94±3.16 70.67±1.76 35.26±3.59 80.10±2.40
AICare 51.37±0.70 85.40±0.48 47.06±1.16 76.23±0.84 49.76±0.86 84.62±0.28 76.07±0.43 45.88±1.12 23.99±2.48 67.35±2.20 42.80±3.79 84.26±2.64
PRISM-${}_{\text{-proto.}}$ 55.52±0.34 87.28±0.11 51.17±0.25 78.66±0.22 55.76±0.90 86.82±0.16 79.12±0.44 50.75±0.65 30.92±2.95 68.87±2.61 50.03±3.97 85.93±1.67
PRISM-${}_{\text{-calib.}}$ 56.16±0.42 87.33±0.22 49.13±2.11 77.87±0.88 55.18±0.77 86.57±0.20 78.66±0.45 50.62±0.70 30.42±2.96 71.07±2.13 46.92±3.20 84.69±1.39
PRISM 57.02±0.38 87.34±0.22 51.31±1.02 78.76±0.59 55.92±0.75 86.82±0.20 79.14±0.33 51.04±0.70 33.60±3.41 73.47±1.11 50.41±3.63 85.82±1.43
6.1 Experimental Results
Table 3 depicts the performance evaluation of baseline methods, PRISM, and its reduced versions for ablation study on four datasets under two prediction tasks. Additionally, we conduct t-test based on the AUPRC metric, the PRISM’s performance improvement against all models are all statistically significant with p-value $<$ 0.01, which underscores that PRISM significantly outperforms all baseline models. Specifically, PRISM outperforms models focused solely on enhancing feature representations with attention mechanisms (like RETAIN, AdaCare, ConCare, AICare, SAFARI), by integrating missing feature status into these mechanisms for improved attention calibration and feature representation. It also exceeds models using similar patient information (such as GRASP, M3Care), showing the value of missing feature status in refining prototype patient representations for better performance. PRISM’s advantage over GRU-D, which only considers local patient visit-based feature missing status, highlights the significance of a global perspective on overall feature missing rates for effective feature representation across patients.
6.2 Ablation Study
Comparing with Reduced Versions
PRISM outperforms PRISM-${}_{\text{-proto.}}$ and PRISM-${}_{\text{-calib.}}$ on main metric AUPRC. This indicates that the two designed learners can enhance patient feature representations from different perspectives: the patient’s individual health data utilized by the feature confidence learner based on the attention mechanism and the prototype similar patient representations utilized by the prototype similar patient learner.
Comparing with Internal Components
To deeply explore the impact of components within each module, we conduct experiments in Table 4, showing PRISM outperforms all baselines. The symbol $\phi$ denotes similarity measure, detailed in Equation 14. The term $\bm{z}$ alone indicates the use of L2 distance for patient similarity measure, whereas $\bm{z},\bm{C}$ additionally incorporates feature confidence, enhancing the model’s discriminative capability. Comparing the roles of various components within the feature confidence learner, the performance when considering both global feature missing rate $\rho$ and local patient’s time interval $\tau$ is higher than considering any single component, which illustrates the necessity of considering the feature missing status from both global and local perspectives. When only considering the local perspective, its performance actually worsens, which is consistent with our observation of the performance of GRU-D. As for confidence-aware prototype patient learner, the performance of confidence-aware patient similarity measurement surpasses that without considering feature confidence, which also shows the impact of missing feature status on measuring similar patients.
Table 4: Performance comparison of internal components on the MIMIC-IV mortality prediction task. “Feat. Conf.” means “Feature Confidence” and “Sim. Meas.” denotes “Similarity Measure”. Bold denotes the best performance within each components, Red denotes the highest performance among all comparisons.
Comparisons Components Metrics
+$\rho$ +$\tau$ +$\phi$ AUPRC($\uparrow$) AUROC($\uparrow$)
Feat. Conf. / / / 54.42±0.43 86.22±0.30
$\surd$ / / 55.11±0.53 86.56±0.33
/ $\surd$ / 52.67±2.52 86.43±0.55
$\surd$ $\surd$ / 55.76±0.90 86.82±0.16
Sim. Meas. / / $\bm{z}$ 55.18±0.77 86.57±0.20
/ / $\bm{z}$, $\bm{C}$ 55.25±0.75 86.60±0.19
PRISM $\surd$ $\surd$ $\bm{z}$, $\bm{C}$ 55.92±0.75 86.82±0.20
6.3 Observations and Analysis
Robustness to Data Sparsity
To assess PRISM’s performance under conditions of data sparsity, we compare it with leading models such as GRASP, MICE(LSTM), and RETAIN in MIMIC-IV in-hospital mortality prediction task. At the volume-level in Figure 2(a), we reduce the data samples in training set, while at the feature-level in Figure 2(b), we intentionally increase the missing feature rates beyond the original missing rate. PRISM excels in both settings. Notably, in situations of extreme data sparsity, such as using only 10% training data and with an overall 97.47% feature missing rate, PRISM significantly outperforms the other models, highlighting its robustness in handling sparse data.
(a) Volume-level Sparsity
(b) Feature-level Sparsity
Figure 2: AUPRC performance across 5 sparsity levels in MIMIC-IV in-hospital mortality prediction task. PRISM significantly outperforms other models in extremely sparse scenarios on both sparsity settings.
Sensitiveness to Cohort Size and Effectiveness of Missing-Feature-Aware Module
We conduct a detailed analysis to examine the impact of prototype patients cohort diversity and the role of a missing-feature-aware module in patient prototypes. Figure 3 shows that integrating a missing-feature-aware module consistently enhances performance across various cluster sizes, as indicated by the superior AUPRC and F1 score. Furthermore, the relatively consistent performance across different cluster sizes demonstrates that our model is not overly sensitive to the number of clusters, highlighting its adaptability and robustness in managing various cohort sizes.
(a) AUPRC
(b) F1
Figure 3: AUPRC and F1 score performance on various prototype patient cohort size in MIMIC-IV in-hospital mortality prediction task. With the missing-feature-aware module, PRISM outperforms its counterpart without. It also shows PRISM is not sensitive to the cohort size.
Variations of Similarity Measures
We conduct a comparison against standard similarity metrics, as detailed in Table 5, PRISM outperforms cosine similarity, and L1 and L2 distances in all metrics. This demonstrates the effectiveness of PRISM’s confidence-aware patient similarity measure in handling sparse EHR data.
Table 5: Comparing different similarity measures in MIMIC-IV in-hospital mortality prediction task.
Measure AUPRC($\uparrow$) AUROC($\uparrow$) F1($\uparrow$)
Cosine 55.58±0.43 86.81±0.10 45.44±1.99
L1 55.59±0.54 86.73±0.20 45.20±1.91
L2 55.38±0.51 86.76±0.09 43.69±1.25
PRISM 55.92±0.75 86.82±0.20 45.47±2.26
Model Efficiency and Complexity
We evaluate the efficiency and complexity against SOTA EHR-specific models baseline models. Table 6 highlights PRISM’s balance between a relatively low total parameter count (215K) and a competitive runtime (69.03s).
Table 6: Computational comparison with latest SOTA EHR-specific models on the MIMIC-IV dataset. All models employ the same 128 hidden dimension and 1024 batch size hyperparameters.
Model # Parameter Runtime (5 epochs)
ConCare (AAAI-2020) 3.3M 96.62s
GRASP (AAAI-2021) 3.3M 97.68s
M3Care (KDD-2022) 469K 68.29s
SAFARI (TKDE-2022) 1.8M 78.97s
AICare (Patterns-2023) 2.2M 72.97s
PRISM 215K 69.03s
Table 7 showcase PRISM’s data preparation time in preprocessing and the overall time for training&validation, and testing on MIMIC datasets. The PRISM pipeline seamlessly computes feature missing statuses in the LOCF pipeline with little extra cost. Moreover, the overall pipeline can be completed in 10 minutes.
Table 7: Data preparation time comparison. Prep. denotes Preprocessing, Val. denotes Validation. All with 50 epochs of training, validate at the end of each epoch. The units are seconds.
Dataset MIMIC-III MIMIC-IV
Prep. w/o Impute LOCF +Ours w/o Impute LOCF +Ours
Time 77.06 300.63 432.19 104.67 358.29 502.68
Pipeline Data Prep. Train+Val. Test Data Prep. Train+Val. Test
Time 432.19 505.83 10.25 502.68 510.73 12.39
Cross-Feature Attention Map
Figure 4 presents the cross-feature attention map from PRISM, contrasting average attention weights with and without feature confidence calibration process. This visualization, based on a single diagnostic record from randomly selected MIMIC-IV patients, plots Key features on the x-axis against Query features on the y-axis. Notably, PRISM reduces attention on features like capillary refill rate, fraction inspired oxygen, and height, which have high missing rates (99.64%, 93.62%, and 99.61%, respectively) in the dataset. PRISM accurately recognizes and calibrates the features with high missing rates, thereby causing the three horizontal lines on the right side of the graph to appear distinctively whiter, showcasing attention-based feature learner module’s interpretability which other baselines lack.
Figure 4: Cross-Feature Attention. Without (Left) / With (Right).
Feature Decay Rates Observation
Figure 5 displays the decay rates of adaptive learning for various features, indicating how their importance diminishes over time. Higher decay rates suggest that the model prioritizes immediate changes in features like heart rate and pH, which is crucial for detecting acute medical conditions such as shock or infection. Conversely, features like sex and systolic blood pressure exhibit lower decay rates, highlighting their relevance in long-term analysis. Notice that the feature ‘Height’ exhibits short-term dynamics, likely due to being absent in over 99% of the data. Consequently, the rapid decay of the ‘Height’ feature, resulting from its high missing rate, does not significantly influence clinical decisions over the long term, which aligns with our intuition.
Figure 5: Feature decay rates observation
7 Conclusions
In this work, we propose PRISM, a prototype patient representation learning framework to address the sparsity issue of EHR data. PRISM perceives and calibrates for missing features, thereby refining patient representations via a confidence-aware prototype patient learner. Significant performance improvements and detailed experimental analysis on four real-world datasets’ in-hospital mortality and 30-day readmission prediction tasks show PRISM’s effectiveness. The work marks a crucial step towards more reliable and effective utilization of EHR data in healthcare, offering a potent solution to the prevalent issue of data sparsity in clinical decision-making.
Ethical Statement
This study, involving the analysis of EHR data is committed to upholding high ethical standards. The employed datasets are all de-identified, ensuring patient confidentiality and privacy.